OpenStudy (anonymous):

Divide these monomials 24abc^4/30a^5bc

OpenStudy (anonymous):

\[\frac{ 24abc ^{4} }{ 30a ^{5}bc }\]

OpenStudy (anonymous):

\[\frac{ 4c^3 }{ 5a^4}\]

OpenStudy (mertsj):

Think of it like this: \[\frac{24}{30}\times\frac{a}{a^5}\times\frac{b}{b}\times\frac{c^4}{c}\]

OpenStudy (anonymous):

@Sky_Pirate

OpenStudy (anonymous):

So what do you do with the 24 and 30?

OpenStudy (mertsj):

Cancel 6 out of both of them: \[\frac{24}{30}=\frac{6\times4}{6\times5}\]